Solve using the elimination method. Show your work. If the system has no solution or an infinite number of solutions, state this.

10x – y = 12.5
3x + y = 7

Answer by mananth(12270)

10 x - y = 12.50 .............1

3 x + y = 7 .............2
Eliminate y


Add the two equations
13 x = 20
/ 13
x = 20/13
plug value of x in (1)
10 x - y = 13
200/13 - y = 13
- y = 13 -200/13
- y = -31/13
y = 31/13
It has a unique solution (20/13, 31/13)